Solve the first order linear inhomogeneous differential equation using the constant variation method

(2x+1)y,=4x+2y


1
Expert's answer
2022-02-07T16:12:51-0500

Given a first order inhomogeneous linear differential equation of the form;

"\displaystyle
y\prime+p(x)y=f(x)\cdots\cdots\cdots\color{red}(1)"

using constant variation method, the general solution is given by;

"\displaystyle
y(x)=v(x)e^{q(x)}+ce^{q(x)}",

where "\displaystyle
v\prime(x)=e^{-q(x)}f(x)", "\displaystyle
q(x)=\int[-p(x)]\ dx", and "c" is and arbitrary constant.


Now, the given DE is of the form;

"\displaystyle
y\prime-\frac{2}{2x+1}y=\frac{4x}{2x+1}"

By comparing the given DE with "\color{red}(1)";

"\displaystyle
q(x)=\int[-p(x)]\ dx=\int\left[\frac{2}{2x+1}\right]\ dx=\ln(2x+1)"

"\displaystyle
v\prime(x)=e^{-q(x)\ dx}\times f(x)=e^{-\ln(2x+1)}\times\frac{4x}{2x+1}=\frac{4x}{(2x+1)^2}"

"\displaystyle
\Rightarrow
v(x)=\int\frac{4x}{(2x+1)^2}\ dx=\int\left(\frac{2}{2x+1}-\frac{2}{(2x+1)^2}\right)\ dx"

"\displaystyle
=\ln(2x+1)+\frac{1}{2x+1}"


Thus, the general solution is;

"\displaystyle
y(x)=v(x)e^{q(x)}+Ae^{q(x)}=\left(\ln(2x+1)+\frac{1}{2x+1}\right)e^{\ln(2x+1)}+ce^{\ln(2x+1)}"

"\displaystyle
=\left(\ln(2x+1)+\frac{1}{2x+1}\right)(2x+1)+c\ln(2x+1)"

"\displaystyle
=(2x+1)\ln(2x+1)+1+c(2x+1)"

"=(2x+1)[c+\ln(2x+1)]+1"